OpenStudy (anonymous):

What are the zeros of f(x)=x^3+4x^2+3x

geerky42 (geerky42):

When you find zeros of f(x), you basically have to solve for x in f(x) = 0

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm x^3+4x^2+3x=0 }\) \(\large\color{blue}{ \rm x(x^2+4x+3)=0 }\) \(\large\color{blue}{ \rm x=0~~~~~~~~or,~~~~~~~(x^2+4x+3)=0 }\) \(\large\color{blue}{ \rm x^2+4x+3=0 }\) \(\large\color{blue}{ \rm (x+3)(x+1)=0 ~~~~~~-->~~~~~x=~~-1,~-3 }\)

OpenStudy (anonymous):

thank you